OpenStudy (anonymous):

What are the values of the function y = –2x – 4 for x = 0, 1, 2, and 3

OpenStudy (lyrae):

Replace x with each of the x-values one at a time and perform the necessary arithmetics. For example: x = 0 => y = -2*0 - 4 = -4

OpenStudy (anonymous):

0, –6, –8, –10 –4, –6, –8, –10 –4, –2, 0, 2 0, 6, 8, 10

OpenStudy (lyrae):

If you insert x = 1 into y = –2x – 4 what does y equal?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

please i need it

OpenStudy (lyrae):

You replace x in the y = –2x – 4 with the x you are given. So if x = 0 then y = –2*0 – 4 = -4 x = 1 then y = -2*1 - 4 = ? and so on. Are following what I'm doing?

OpenStudy (anonymous):

-2?

OpenStudy (lyrae):

Is -2 - 4 = -2?

OpenStudy (anonymous):

c?

OpenStudy (lyrae):

Nah -2 - 4 = -6

OpenStudy (anonymous):

b?

OpenStudy (lyrae):

yes :)

OpenStudy (anonymous):

ok thank you :)